mikewwe13:

What is the value of h when the function is converted to vertex form? Note: Vertex form is p(x) = a(x − h)^2 + k . p(x) = x^2 − 14x + 29 Enter your answer in the box. h = ______

Vocaloid:

first we need to re-write the polynomial as a perfect square trinomial take the b-coefficient (remember, the b coefficient is the number being multiplied to x), divide by 2, square the result, let me know what you get

mikewwe13:

The first term is, x2 its coefficient is 1 . The middle term is, -14x its coefficient is -14 . The last term, "the constant", is +29

mikewwe13:

AM I CORRECT ?

Vocaloid:

please follow my instructions

mikewwe13:

14x divided by 2

Vocaloid:

just the coefficient -14, divide that by 2, then square it.

mikewwe13:

-7

Vocaloid:

after you divide it by 2, you must square it what is (-7)^2 =?

mikewwe13:

49

Vocaloid:

good, so we add 49 to the polynomial x^2 − 14x + 49 + 29 however, we must also subtract 49 from the end to keep the polynomial the same x^2 − 14x + 49 + 29 - 49 now, we put this part in parentheses to make our trinomial (x^2 − 14x + 49) + 29 - 49

Vocaloid:

now, factor this part: (x^2 − 14x + 49)

mikewwe13:

(x - 7)^2

Vocaloid:

awesome, then we re-write the polynomial like so: (x-7)^2 + 29 - 49 simplify the 29 - 49 part gives us (x-7)^2 - 20 now, it's in vertex form now let's put the two equations side by side (x-h)^2 + k = (x-7)^2 - 20 what is the value of h?

mikewwe13:

14

Vocaloid:

let's just consider the exponential parts put (x-h)^2 and (x-7)^2 side by side compare the two, what do you think h might be equal to?

mikewwe13:

h^2 - 2hx + x^2

Vocaloid:



Vocaloid:

if x - h = x - 7 what does h equal?

mikewwe13:

h = x - 7

Vocaloid:

x - h = x - 7 subtract x from both sides, what do you get?

mikewwe13:

h = 7

Vocaloid:

awesome so h = 7 = your answer

mikewwe13:

What are the x-intercepts of the quadratic function? f(x) = x^2 − 3x − 10 Enter your answers in the boxes. ______ and ______

Vocaloid:

start by factoring it